if the area of a rectangle is 48 and the length is 4 less than twice the width,what is the measure of the width?
Answer by checkley77(12844)

A=LW
L=2W-4
48=(2W-4)W
48=2W^2-4W
2W^2-4W-48=0
2(W^2-2W-24)=0
2(W-6)(W+4)=0
W-6=0
W=6 ANS.
L=2*6-4=12-4=8 ANS.
PROOF:
48=(2*6-4)6
48=(12-4)6
48=8*6
48=48